what magnetic flux penetrates the flat surface 50. when the field is induced 0.4 t, if this surface: a) is perpendicular to the vector and the field's field; b) is located at an angle of 45 ° to the induction vector; c) is located at an angle of 30 ° to the induction vector
1
Expert's answer
2019-06-24T09:22:24-0400

a)


"\Phi=BS \sin{90}= (0.4)(50)\sin{90}=20\ Wb"

b)


"\Phi=BS \sin{45}= (0.4)(50)\sin{45}=14.1\ Wb"

c)


"\Phi=BS \sin{30}= (0.4)(50)\sin{30}=10\ Wb"